If a pool can hold 2000 litres of water, what percentage is lost when 25 litres evaporates?
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

If a pool can hold 2000 litres of water, what percentage is lost when 25 litres evaporates?
-----
Solve::
x/100 = 25/2000
-----
x = 2500/2000 = 1.25
Answer:: 1.25%
------------------
cheers,
Stan H.

Answer by rfer(16322)

25/2000=0.0125=1.25%